A stock just paid a dividend of $2.14. The dividend is expected to grow at 20.29% for three years and then grow at 4.33% thereafter. The required return on the stock is 11.91%. What is the value of the stock?

The risk-free rate is 1.05% and the market risk premium is 6.41%. A stock with a β of 0.93 just paid a dividend of $2.52. The dividend is expected to grow at 21.12% for three years and then grow at 4.61% forever. What is the value of the stock?

Derek plans to buy a $28,129.00 car. The dealership offers zero percent financing for 59.00 months with the first payment due at signing (today). Derek would be willing to pay for the car in full today if the dealership offers him $____ cash back. He can borrow money from his bank at an interest rate of 4.98%.
Answer #1

Hi

As per policy we will solve only first question here.

Here dividend just paid D0= $2.14

D1 = 2.14*(1+20.29%)

=$2.57

D2 = 2.57*(1+20.29%)

=$3.0965

D3 = 3.0965*(1+20.29%) = $3.72

terminal value = D3*(1+g)/(r-g)

=3.72*(1+4.33%)/(11.91%-4.33%)

=$51.26752

A stock's current value will be sum of present value of its future dividends.

Value of stock = 2.57/(1+11.91%) + 3.0965/(1+11.91%)^2 + 3.72/(1+11.91%)^3 + 51.26752/(1+11.91%)^3

=2.3 + 2.47 + 2.66+36.58

= $44.01

Thanks